OpenStudy (anonymous):

i am practising here ......\(\trying~to~do~latex\)if u want to tell me more then it would be good.........

OpenStudy (anonymous):

\(\huge\color{purple}{\heartsuit}\) \(\huge\color{red}{\heartsuit}\) \(\color{green}{\smile}\) \(\scr{this~is~heart~smile}\)

OpenStudy (anonymous):

\(\squrt{x+3}\) @tester97 how to do this math stuff?

OpenStudy (abhisar):

\(\color{blue}{\text{Originally Posted by}}\) @ziqbal103 \(\huge\ ♥ \ ♥ \) \(\huge~\smile\) \(\scr{this~is~heart~smile}\) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (abhisar):

\(\color{green}{\bigstar}\color{red}{\bigstar} \\ \huge \smile\)

OpenStudy (anonymous):

(\(\sqrt{2})^2\) \(\ (\ sqrt {2 }^2

OpenStudy (anonymous):

\(\huge\color{brown}{how~i~write~degree~and~theta~in~latex}\) @thomaster

OpenStudy (anonymous):

@e.mccormick can you help me please

OpenStudy (anonymous):

\(\longrighttarrow\)) how to write longrighttarrow?

OpenStudy (anonymous):

\(\huge{longrighttarrrow}\) @dls

OpenStudy (anonymous):

\(\huge\long~right~tarrow\)

OpenStudy (e.mccormick):

``` \(\longrightarrow\) ``` \(\longrightarrow\)

OpenStudy (anonymous):

this not

OpenStudy (anonymous):

 this

OpenStudy (anonymous):

\(\huge\implies\) \( \ huge \implies \) \implies\)

OpenStudy (anonymous):

\(\huge\theta\) \(\ huge \theta \ )

OpenStudy (e.mccormick):

That is not a rightarrow, it is a Rightarrow. The capital at the start. \(\Rightarrow\) \(\Longrightarrow\)

OpenStudy (anonymous):

\(\huge=\) \ ( \huge \=\ )

OpenStudy (e.mccormick):

\(\rightarrow\) \(\leftarrow\) \(\leftrightarrow\) \(\longrightarrow\) \(\longleftarrow\) \(\longleftrightarrow\) \(\Rightarrow\) \(\Leftarrow\) \(\Leftrightarrow\) \(\Longrightarrow\) \(\Longleftarrow\) \(\Longleftrightarrow\) \(\Rrightarrow\) Bla bla bla. Here: http://www.access2science.com/latex/Arrows.html

OpenStudy (anonymous):

\(45^o\) \(\huge45^0\) \(\Huge45^0\) ( 45^0\)

OpenStudy (anonymous):

\(\sqrt7{3+5}\) \(\F(x)\) \[\Huge\cancel{f(x)}\] \[ Hugecancel {f)(x)\}

OpenStudy (anonymous):

\(\Huge\(\color{green}{I~am~practising~what~about~you?}\) \(\Huge\color{red}\LaTeX\triangle\) \(\Huge\pm56 ............\) \(\Huge99\approx100......\) \(\Huge0\leq8\) \(\Huge\longleftarrow\) \leq) \(\Huge\LaTeX\triangle\) \(\Huge\in\) \(\longleftarrow\) \Huge\in\) \(\huge\leftrightarrow\) \(\Huge\10\geq2\) \(\geq)

OpenStudy (anonymous):

\(\Huge\color{orange}{Hey~what~are~you~looking~here?@:( }\)